# Adaptive Partition Factor Analysis

### v2

@@ Abstract
Factor Analysis has traditionally been utilized across diverse disciplines to extrapolate latent traits that influence the behavior of multivariate observed variables. Historically, the focus has been on analyzing data from a single study, neglecting the potential study-specific variations present in data from multiple studies. Multi-study factor analysis has emerged as a recent methodological advancement that addresses this gap by distinguishing between latent traits shared across studies and study-specific components arising from artifactual or population-specific sources of variation. In this paper, we extend the current methodologies by introducing novel shrinkage priors for the latent factors, thereby accommodating a broader spectrum of scenarios—from the absence of study-specific latent factors to models in which factors pertain only to small subgroups nested within or shared between the studies. For the proposed construction we provide conditions for identifiability of factor loadings and guidelines to perform straightforward posterior computation via Gibbs sampling. Through comprehensive simulation studies, we demonstrate that our proposed method exhibits competing performance across a variety of scenarios compared to existing methods, yet providing richer insights. The practical benefits of our approach are further illustrated through applications to bird species co-occurrence data and ovarian cancer gene expression data.
Keywords: Factor models, Neural networks, Shrinkage priors, Sparsity.
@@ 1 Introduction
In these contexts, factor analysis is a standard technique for identifying common patterns of variation in multivariate data and relating it to hidden causes, the so-called latent factors. However, a key challenge arises when the latent factors identified across different studies combine both shared factors (common across studies) and factors unique to individual studies. This issue is particularly pronounced in our motivating contexts of animal co-occurrence studies and genomics. For instance, species occurrence data collected across spatial-temporal surveys often exhibit site-specific deviations despite shared environmental conditions (Ovaskainen et al., 2017). Similarly, high-throughput genomic experiments are prone to technical variability, where artifacts from experimental conditions may dominate biological signals (Irizarry et al., 2003; Shi et al., 2006; Vallejos et al., 2017; Hicks et al., 2018). While preprocessing adjusts for some biases, residual heterogeneity in covariance structures often persists across groups, challenging assumptions of homogeneity (Avalos-Pacheco et al., 2022).
@@
where $\epsilon_{is}$ represents a zero-mean idiosyncratic error term and $\eta_{is}$ is a $d$-dimensional set of shared latent factors, with $d\ll p$ and with $\Lambda$ the corresponding factor loading matrix of dimension $p\times d$. Similarly, $\Gamma_{s}$ is a study-specific factor loading matrix of dimension $p\times k_{s}$, with $k_{s}\ll p$, and ${\varphi}_{is}$ its corresponding latent factors. Notably, each subject belongs to a single study and, in fact, the index $s$ in the notation $y_{is}$ may be redundant as also discussed later in Section 2. Figure 1 reports a pictorial representation of the data matrix $Y$ with $n$ rows defined as $Y=(y_{i1}^{\top},\dots,y_{nS}^{\top})^{\top}$.
@@
In all of these approaches there is however a modeling limitation, i.e. they allow precisely $S$ study-specific loading matrices $\Gamma_{s}$. In practice, this assumption is often violated. We provide some examples, though a continuous spectrum of possibilities exists: (i) two or more studies exhibit high homogeneity and may share identical or nearly identical latent representations, and (ii) one study might contain highly heterogeneous subject groups, potentially forming distinct sub-populations with different latent structures (as formalized in Equation (1)). This could arise from minor experimental variations or unmeasured confounders. Additional possibilities include (iii) the existence of a subset of units from different groups that share common characteristics, or (iv) instances where some units might simultaneously exhibit characteristics of multiple groups due to various reasons.
@@
The next section details APAFA in terms of model specification, prior distributions, and provides practical guidelines for prior elicitation. We also explore the connections to neural networks and present generalizations. In addition, we address model identifiability. Section 3 reports a comprehensive empirical assessment through simulation studies covering various data-generating scenarios. Section 4 reports two illustrative data analyses further highlighting the advantages of our approach in the motivating examples of bird species occurrence and genomics. The final discussion section explores generalizations and potential extensions. The code to reproduce our analyses available at https://github.com/elenabortolato/APAFA. The Supplementary Materials provide additional technical details on posterior computation methods, extended empirical results with further experimental analyses, and implementation specifications.
@@ 2.1 Model and prior specification
where $\Gamma=(\Gamma_{1},\dots,\Gamma_{S})$ concatenates along columns all the study-specific factor loading matrices into a $p\times k$ matrix with $k=\sum_{s=1}^{S}k_{s}$ and ${\varphi}_{i}$ is a $k$-dimensional augmented vector containing the ${\varphi}_{is}$ in Equation (1) framed with suitable pattern of zeros. Specifically, all the latent factors pertaining to group $s$ will have $\sum_{l=1}^{s-1}k_{l}$ zeros, followed by subject-specific ${\varphi}_{is}$, followed by $\sum_{l=s+1}^{S}k_{l}$ zeros. This formulation introduces two distinct factor loading matrices: $\Lambda$ containing loadings common to all units, and $\Gamma$ containing all the study-specific loadings that are activated following the non-zero patterns of $\varphi_{i}$. An illustration is depicted in Figure 1.
$\mathbf{Y}$$\mathbf{H}$$\mathbf{\Lambda^{\top}}$$\mathbf{\Phi}$$\mathbf{\Gamma}^{\top}$$\mathbf{\epsilon}$===++++++
@@
Under this characterization, one can think to have, for each $i$, a vector of dummy variables $x_{i}$ with $S$ entries characterizing the study to which unit $i$ belongs to and then assume that the $h$-th element of the vector $\varphi_{i}$ is $\varphi_{ih}=\tilde{\varphi}_{ih}\psi_{ih}$ where $\tilde{\varphi}_{ih}$ is a continuous random variable and $\psi_{ih}=f_{h}(x_{i})$ with $f_{h}$ a deterministic activation function. In De Vito et al. (2019, 2021), for example, $f_{h}(x_{i})=x_{i}^{\top}1_{S}$ where $1_{S}$ is a $S$-dimensional vector of ones. In this paper, we propose to incorporate the information contained in the $x_{i}$’s in a more flexible manner letting
@@
Consistently with these approaches, but focusing on our motivating factor analysis context, we now define the prior structure for model (2). Specifically, we let
@@
where the variance of $\varphi_{ih}$ is a product of a global scale dependent on the index $h$ and a local scale dependent on the indexes $i$ and $h$, and most importantly, on the variable $x_{i}$ through Equation (4).
We let the global scale parameter $\tau^{\varphi}_{h}\sim\text{Ber}(1-\rho_{h})$, where $\{\rho_{h}\}\sim\mbox{CUSP}(\alpha^{\varphi})$ follows a cumulative shrinkage process (Legramanti et al., 2020), i.e. for $h=1,2,\dots,$ $\rho_{h}=\sum_{l\leq h}w^{\varphi}_{l}$,   $w^{\varphi}_{l}=v^{\varphi}_{l}\prod_{m<l}(1-v^{\varphi}_{l})$, and $v^{\varphi}_{l}\sim\text{Beta}(1,\alpha^{\varphi})$. This approach facilitates learning the number of study-specific factors or, following the neural network interpretation of our model, the number of neurons in the hidden layer. Notably, this construction, inspired by successful applications in factor models (Legramanti et al., 2020; Schiavon et al., 2022), offers valuable insights into performing Bayesian inference for the number of nodes in the hidden layer of a Bayesian neural network, potentially connecting with existing literature on this topic (e.g Wen et al., 2016; Nalisnick et al., 2019; Fortuin, 2022; Cui et al., 2022; Jantre et al., 2023; Sell & Singh, 2023).
We now move the discussion on the local scales $\psi_{ih}(x_{i})$ that are assumed to be Bernoulli variables with probability depending on $x_{i}$. Specifically, we let
$\psi_{ih}(x_{i})\sim\text{Ber}\{\text{logit}^{-1}(x_{i}^{\top}\beta_{h})\}.$ (6)
Note that under this specification, we can also write $\varphi_{ih}(x_{i})=\tilde{\varphi}_{ih}\psi_{ih}(x_{i})$ and
@@
where $\delta_{a}$ is a Dirac mass at value $a$. Conversely, marginalizing out $\tilde{\varphi}_{ih}$ under unit variance leads to
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}y_{i}\sim N(0,\Omega_{i}),\quad\Omega_{i}=\Lambda\Lambda^{\top}+\Gamma\text{diag}{\{\psi_{i}\}}\Gamma^{\top}+\Sigma.}$ (7)
Notably (6) corresponds to a simple form of conditional variational autoencoder that, differently from the standard, allows for the latent covariance structure to be informed by the independent variables. Variational autoencoders are a specific extension of neural networks, widely recognized as a descendant of classical factor models (Kingma & Welling, 2014; Sen et al., 2024).
@@
We complete the prior specification for the study-specific part assuming standard prior for the factor loadings. Specifically for the general element $\gamma_{jh}$ of the factor loading matrix $\Gamma$, we let $\gamma_{jh}\sim N(0,\zeta^{\gamma}_{h})$, with $\zeta^{\gamma}_{h}\sim\text{IGa}(a_{\gamma},b_{\gamma}).$
One remark on the activation function pertains to the potential lack of properness of the conditional posterior distribution for the parameters $\beta_{h}$, $h=1,\dots,k$. This concern is tied to the issue of perfect separation in logistic regression. Assuming the group structure encoded by the $x_{i}$ dummy variables is both accurate and relevant to the problem at hand—in other words, that the original MSFA model by De Vito et al. (2021) is correctly specified— a perfect separation in the $\psi_{ih}$ occurs with respect to the $x_{i}$ within the logistic regression. For this reason, the prior distribution on each parameter $\beta_{h}$ should be sufficiently concentrated around zero to prevent the associated conditional posterior from exhibiting a tendency toward monotonicity. To this end, we specify $\beta_{h}\sim N(0,B)$, where $B$ is a $S\times S$ diagonal matrix with elements proportional to $n^{-1}$. Thus, when perfect separation in the $\psi_{ih}$’s with respect to the $x_{is}$’s occurs, the full conditional posterior is
@@
from which we can also recognize that if the number of groups $S$ increases the posterior concentrates close to 0, for any group size. Thus, despite the apparent inconsistency of a prior that induces skepticism regarding the given group configuration, it facilitates posterior inference.
@@
Given the model, posterior computation is performed using Gibbs sampling, with full methodological details provided in the Supplementary Materials.
2.2 Model properties
Identifiability of factor models is an overarching issue. While these models are appealing in revealing possible interpretable relationships between a small set of latent factors and multivariate observations—even with complex dependency structures—this interpretability comes at a cost: factor analytic models lack identifiability in fundamental ways. For instance, loadings are not identifiable due to rotation, sign and permutation invariance. Rotational invariance refers to the transformation of loadings and factors by an orthogonal matrix, sign switching arises because one can simultaneously change the signs of certain elements without affecting the overall structure, while permutation invariance occurs because there is no intrinsic ordering in the factor loading matrix columns. See Papastamoulis & Ntzoufras (2022) Xu et al. (2023) Frühwirth-Schnatter et al. (2024) for a comprehensive review.
Identification issues are exacerbated in the MSFA framework of De Vito et al. (2021) where it may be difficult to disentangle the signal between shared and study-specific factors, a problem known as information switching. For instance, Chandra et al. (2024) shows how a linear combination of study-specific factors can instead be represented and reincorporated as part of the shared component. Building on these concepts, we provide conditions to identify the shared from the specific factors in our setup. Throughout, let $\mathcal{O}_{k}$ denote the space of $k\times k$ orthogonal matrices. We formalize the notion of information switching through the following definition:
@@ 2.2 Model properties
Since $\text{rank}(G_{s})=k$, then $\text{rank}(G)=S_{n}k$, and $\text{rank}(G)=\text{rank}(G|\mathbf{w})$ by construction. Thus, since $k<p(p+1)/2$ by assumption, the solution is unique by the Rouché–Capelli Theorem. The uniqueness of the solution for the elements $\psi_{s}^{*}$ ensures that it is impossible to find a representation where for one $h$, $\tilde{\Psi}_{h}$ is a vector of one.
The proof of Theorem 1 is based on proving existence of unique vectors $\psi_{s}^{*}$ for the specific factors. This guarantees that no alternative representation exists in which, for some $h$, the corresponding $\tilde{\Psi}_{h}$ is a vector of ones. Notably, it also establishes the uniqueness of the sparsity patterns encoded by the $\Psi$ matrix. That is, any alternative representation inducing the same covariance structure must preserve the subject-specific activation patterns of the specific components. The following corollary formalizes this consequence.
@@
Based on the identifiability findings of Theorem 1, the practical conditions that ensures resistance to information switching in APAFA are discussed in Lemma 1 and 2.
@@
Under any continuous prior for the elements of $\Gamma$ and truncating its number of columns to $K=p(p+1)/2-1$, the information switching for model (7) has zero probability.
@@
From Markov’s inequality $\text{pr}(k>2(p+1)/2)\leq\frac{\alpha^{\varphi}}{p(p+1)/2}\leq\varepsilon$
A key advantage of our model is that the requirement of $k<p(p+1)/2$ is minimal as it is customary in factor analysis to have latent factors of a considerably lower dimension than the observed data ($k<p$). Notably, Chandra et al. (2024), in different settings, necessitate the dimension of the specific part to be smaller than that of the shared part implicitly considering the study-specific part as a noise or disturbance parameter. Although we are in different settings, where the number and composition of the groups is not pre-specified, we do not restrict $k$ to be also smaller than $d$. This is crucial in many contexts. For instance, in cancer studies, it is possible to study very different types of cancer simultaneously, identifying a range of specific factors that describe the specific characteristics of each type of cancer, while also uncovering a few common factors that are shared across cancers, regardless of the organ or tissue affected possibly pertaining to fundamental mechanisms of oncogenesis.
@@
We begin by fixing the ordering of the columns of $\Gamma$. Under this fixed ordering, we will show that the only admissible orthogonal matrices $P\in\mathcal{O}_{k}$ are diagonal with entries ±$\pm 1$. Notably, applying any permutation to the columns of $\Gamma$ initially and repeating the argument would yield the same conclusion, i.e. the only admissible $P$ relative to the permuted ordering are diagonal with entries ±$\pm 1$. Consequently, the general solution requires $P$ to be a signed permutation matrix.
@@
and define $\Gamma_{s}\in\mathbb{R}^{p\times k_{s}}$ to be the submatrix of $\Gamma$ consisting of the columns indexed by $A_{s}$. Notably these are identifiable thanks to Corollary 1.
The proof proceeds considering that $P_{s}\in\mathcal{O}_{k_{s}}$, leaves $\Gamma^{\prime}_{s}\Gamma^{\prime}_{s}{}^{\top}$ invariant and hence that the marginal distribution of group $s$ is also invariant under such rotations. Given these $P_{s}\in\mathcal{O}_{k_{s}}$ for each $s=1,\dots,S_{n}$, we investigate whether a “global” orthogonal matrix $P\in\mathcal{O}_{k}$ can be constructed by assembling the blocks $P_{s}$. This requires consistency across overlapping blocks, ensuring that each $P_{s}$ remains orthogonal, for example satisfying $\sum_{l\in A_{s}}\|P_{s,hl}\|=1$ for all $h$ and $s$. We will argue in the following that, under the Non-replicable Sparsity Pattern Condition of Definition 2, this is possible only if $P$ is a diagonal matrix with entries ±$\pm 1$.
For any $h=1,\dots,k$ there exist at least one group loading factor $h$, i.e. $\psi^{*}_{sh}=1$ for at least one $s$. Then we have two possibilities: $i)$ $\psi_{s^{\prime}h}^{*}=0$ for all other $s^{\prime}\neq s$, or $ii)$ $\psi_{s^{\prime}h}^{*}=1$ for at least another $s^{\prime}\neq s$.
In the first case $i)$, $h$ is factor belonging exclusively to group $s$. Notably, by Definition 2 it is the only exclusive factor. If this is the only factor of group $s$, then the rotation matrix $P_{s}$ reduces to a scalar. Since $P$ is orthogonal, $\|P_{h}\|^{2}=1$ implies $P_{hh}=\pm 1$ and all other entries in the $h$th row/column of $P$ must vanish. If $h$ is not the only factor of group $s$, then $A_{s}$ contains at least another $h^{\prime}$ that is shared among $s$ and at least another $s^{\prime}$. In this case, $h^{\prime}$ is in case $ii)$. Consider then case $ii)$. In order to be orthogonal, matrix $P$ needs to satisfy
 $\>1=\|P_{h}\|^{2}=\sum_{l=1}^{k}P_{lh}^{2}=\sum_{l\in A_{s^{\prime}}\cap A_{s}}P_{lh}^{2}+\sum_{l\in A_{s^{\prime}}\setminus A_{s}}P_{lh}^{2}+\sum_{l\in A_{s}\setminus A_{s^{\prime}}}P_{lh}^{2}+\sum_{l\notin A_{s}\cup A_{s^{\prime}}}P_{lh}^{2}.$ (8)
@@
 $\>1=\|P_{{s},h}\|^{2}=\sum_{l\in A_{s}\cap A_{s^{\prime}}}P_{lh}^{2}+\sum_{l\in A_{s}\setminus A_{s^{\prime}}}P_{lh}^{2}$ (9)

### v3

@@ Abstract
Factor Analysis has traditionally been utilized across diverse disciplines to extrapolate latent traits that influence the behavior of multivariate observed variables. Historically, the focus has been on analyzing data from a single study, neglecting the potential study-specific variations present in data from multiple studies. Multi-study factor analysis has emerged as a recent methodological advancement that addresses this gap by distinguishing between latent traits shared across studies and study-specific components arising from artifactual or population-specific sources of variation. In this paper, we extend the current Bayesian methodologies by introducing novel shrinkage priors for the latent factors, thereby accommodating a broader spectrum of scenarios—from the absence of study-specific latent factors to models in which factors pertain only to small subgroups nested within or shared between the studies. For the proposed construction we provide conditions for identifiability of factor loadings and guidelines to perform straightforward posterior computation via Gibbs sampling. Through comprehensive simulation studies, we demonstrate that our proposed method exhibits competing performance across a variety of scenarios compared to existing methods, yet providing richer insights. The practical benefits of our approach are further illustrated through applications to bird species co-occurrence data and ovarian cancer gene expression data.
Keywords: Bayesian inference, Factor models, Neural networks, Shrinkage priors, Sparsity.
@@ 1 Introduction
In these contexts, factor analysis is a standard technique for identifying common patterns of variation in multivariate data and relating it to hidden causes, the so-called latent factors. However, a key challenge arises when the latent factors identified across different studies combine both shared factors (common across studies) and factors unique to individual studies. This issue is particularly pronounced in our motivating contexts of animal co-occurrence studies and genomics. For instance, species occurrence data collected across spatial-temporal surveys often exhibit site-specific deviations despite shared environmental conditions (Ovaskainen et al., 2017). Similarly, high-throughput genomic experiments are prone to technical variability, where artifacts from experimental conditions may dominate biological signals (Irizarry et al., 2003; Shi et al., 2006; Vallejos et al., 2017; Hicks et al., 2018). While preprocessing adjusts for some biases, residual heterogeneity in covariance structures often persists across groups, challenging homogeneity assumption (Avalos-Pacheco et al., 2022).
@@
where $\epsilon_{is}$ represents a zero-mean idiosyncratic error term and $\eta_{is}$ is a $d$-dimensional set of shared latent factors, with $d\ll p$ and with $\Lambda$ the corresponding factor loading matrix of dimension $p\times d$. Similarly, $\Gamma_{s}$ is a study-specific factor loading matrix of dimension $p\times k_{s}$, with $k_{s}\ll p$, and ${\varphi}_{is}$ its corresponding latent factors. Notably, each subject belongs to a single study and, in fact, the index $s$ in the notation $y_{is}$ may be redundant as also discussed later in Section 2. Figure 1 reports a pictorial representation of the data matrix $Y$ with $n$ rows defined as $Y=(y_{11}^{\top},\dots,y_{n_{S}S}^{\top})^{\top}$.
@@
In all of these approaches there is however a modeling limitation, i.e. they allow precisely $S$ study-specific loading matrices $\Gamma_{s}$. In practice, this assumption is often violated. We provide some examples, though a continuous spectrum of possibilities exists: (i) two or more studies exhibit high homogeneity and may share identical or nearly identical latent representations, and (ii) one study might contain highly heterogeneous subject groups, potentially forming distinct sub-populations with different latent structures. This could arise from minor experimental variations or unmeasured confounders. Additional possibilities include (iii) the existence of a subset of units from different groups that share common characteristics, or (iv) instances where some units might simultaneously exhibit characteristics of multiple groups due to various reasons.
@@
The next section details APAFA in terms of model specification, prior distributions, and provides practical guidelines for prior elicitation. We also explore the connections to neural networks and present generalizations. In addition, we address model identifiability. Section 3 reports a comprehensive empirical assessment through simulation studies covering various data-generating scenarios. Section 4 reports two illustrative data analyses further highlighting the advantages of our approach in the motivating examples of bird species occurrence and genomics. The final discussion section explores generalizations and potential extensions. The code to reproduce our analyses is available at https://github.com/elenabortolato/APAFA.
@@ 2.1 Model specification
where $\Gamma=(\Gamma_{1},\dots,\Gamma_{S})$ concatenates along columns all the study-specific factor loading matrices into a $p\times k$ matrix with $k=\sum_{s=1}^{S}k_{s}$ and ${\varphi}_{i}$ is a $k$-dimensional augmented vector containing the ${\varphi}_{is}$ in Equation (1) framed with suitable pattern of zeros. Specifically, all the latent factors pertaining to group $s$ will have $\sum_{l=1}^{s-1}k_{l}$ zeros, followed by subject-specific ${\varphi}_{is}$, followed by $\sum_{l=s+1}^{S}k_{l}$ zeros. This formulation introduces two distinct factor loading matrices: $\Lambda$ containing loadings common to all units, and $\Gamma$ containing all the study-specific loadings that are activated following the non-zero patterns of $\varphi_{i}$. An illustration is depicted in Figure 1 for $n=15$, $p=5$, $d=2$, $k=4$, $S=3$, $k_{1}=k_{2}=1$, and $k_{3}=2$.
$\mathbf{Y}$$\mathbf{H}$$\mathbf{\Lambda^{\top}}$$\mathbf{\Phi}$$\mathbf{\Gamma}^{\top}$$\mathbf{\epsilon}$==++++
@@
Under this characterization, one can think to have, for each $i$, a vector of dummy variables $x_{i}$ with $S$ entries characterizing the study to which unit $i$ belongs to and then assume that the $h$-th element of the vector $\varphi_{i}$ is $\varphi_{ih}=\tilde{\varphi}_{ih}\psi_{ih}$ where $\tilde{\varphi}_{ih}$ is a continuous random variable and $\psi_{ih}=f_{h}(x_{i})$ with $f_{h}$ a deterministic activation function. In De Vito et al. (2019, 2021), for example, $f_{h}(x_{i})=x_{i}^{\top}1_{S}$ where $1_{S}$ is a $S$-dimensional vector of ones. In this paper, we propose to incorporate the information contained in the $x_{i}$’s in a more flexible manner, i.e.
@@
2.2 Prior specification
We now define the prior structure for model (2). Specifically, we let
@@ 2.2 Prior specification
where the variance of $\varphi_{ih}$ is a product of a global scale (depending on the index $h$) and a local scale (dependent on both indexes $i$ and $h$), and most importantly, on the variable $x_{i}$ through Equation (4).
We let the global scale parameter $\tau^{\varphi}_{h}\sim\text{Ber}(1-\rho_{h})$, where $\{\rho_{h}\}\sim\mbox{CUSP}(\alpha^{\varphi})$ follows a cumulative shrinkage process (Legramanti et al., 2020), i.e. for $h=1,2,\dots,$ $\rho_{h}=\sum_{l\leq h}w^{\varphi}_{l}$,   $w^{\varphi}_{l}=v^{\varphi}_{l}\prod_{m<l}(1-v^{\varphi}_{m})$, and $v^{\varphi}_{l}\sim\text{Beta}(1,\alpha^{\varphi})$. This approach facilitates learning the number of study-specific factors or, following the neural network interpretation of our model, the number of neurons in the hidden layer. This construction, inspired by successful applications in factor models (Legramanti et al., 2020; Schiavon et al., 2022), offers valuable insights into performing Bayesian inference for the number of nodes in the hidden layer of a Bayesian neural network, potentially connecting with existing literature on this topic (e.g Wen et al., 2016; Nalisnick et al., 2019; Fortuin, 2022; Cui et al., 2022; Jantre et al., 2023; Sell & Singh, 2023).
We now move the discussion to the local scales $\psi_{ih}(x_{i})$ that are assumed to be Bernoulli variables with probability depending on $x_{i}$. Specifically, we let
$\psi_{ih}(x_{i})\sim\text{Ber}\{f_{h}(x_{i}^{\top}\beta_{h})\},$ (6)
with $f_{h}(\cdot)=\text{logit}^{-1}(\cdot),$ for all $h$. Since $\varphi_{ih}(x_{i})=\tilde{\varphi}_{ih}\psi_{ih}(x_{i})$, Equations (5) and (6) imply
@@
where $\delta_{a}$ is a Dirac mass at value $a$. Conversely, marginalizing out $\tilde{\varphi}_{ih}$ while keeping $\varphi_{i}$ fixed and $\tau_{h}^{\psi}=1$ leads to
$y_{i}\sim N(0,\Omega_{i}),\quad\Omega_{i}=\Lambda\Lambda^{\top}+\Gamma\text{diag}{\{\psi_{i}\}}\Gamma^{\top}+\Sigma.$} (7)
Notably, (6) corresponds to a simple form of a conditional variational autoencoder that, unlike the standard version, allows the latent covariance structure to be informed by the independent variables. Variational autoencoders are a specific extension of neural networks, widely recognized as descendants of classical factor models (Kingma & Welling, 2014; Sen et al., 2024).
@@
We complete the prior specification for the specific part assuming standard prior for the factor loadings. Specifically for the general element $\gamma_{jh}$ of the factor loading matrix $\Gamma$, we let $\gamma_{jh}\sim N(0,\zeta^{\gamma}_{h})$, with $\zeta^{\gamma}_{h}\sim\text{IGa}(a_{\gamma},b_{\gamma}).$
One remark on the activation function pertains to the potential lack of properness of the conditional posterior distribution for the parameters $\beta_{h}$, $h=1,\dots,k$. This concern is tied to the issue of perfect separation in logistic regression. Assuming the group structure encoded by the $x_{i}$ dummy variables is both accurate and relevant to the problem at hand—in other words, that the original MSFA model by De Vito et al. (2021) is correctly specified— a perfect separation in the $\psi_{ih}$ may occur with respect to the $x_{i}$ within the logistic regression. For this reason, the prior distribution on each parameter $\beta_{h}$ should be sufficiently concentrated around zero to prevent the associated conditional posterior from exhibiting a tendency toward monotonicity. To this end, we specify $\beta_{h}\sim N(0,B)$, where $B$ is a $S\times S$ diagonal matrix with elements proportional to $n^{-1}$. Thus, when perfect separation in the $\psi_{ih}$’s with respect to the $x_{i}$’s occurs, the full conditional posterior is
@@
from which we can also recognize that if the number of groups $S$ increases the posterior concentrates close to 0, for any group size. Despite the apparent inconsistency of a prior inducing skepticism on the group configuration, it facilitates posterior inference.
@@
Posterior inference is conducted via Gibbs sampling. See the Supplementary Material.
2.3 Identification properties
Identifiability of factor models is an overarching issue. While these models are appealing in revealing possible interpretable relationships between a small set of latent factors and multivariate observations—even with complex dependency structures—this interpretability comes at a cost: factor analytic models lack identifiability in fundamental ways. For instance, loadings are not identifiable due to rotation, sign and permutation invariance. Rotational invariance allows transformation of loadings and factors by an orthogonal matrix, sign invariance arises when simultaneous sign changes leave the structure unchanged, and permutation invariance reflects the lack of intrinsic ordering in loading matrix columns. See Papastamoulis & Ntzoufras (2022) Xu et al. (2023) Frühwirth-Schnatter et al. (2025) for a comprehensive review.
Identification issues are exacerbated in the MSFA framework of De Vito et al. (2021) where it may be difficult to disentangle the signal between shared and specific factors, a problem known as information switching. For instance, Chandra et al. (2025) show how a linear combination of study-specific factors can instead be represented and reincorporated as part of the shared component. Building on these concepts, we provide conditions to identify the shared from the specific factors in our setup. Throughout, let $\mathcal{O}_{k}$ denote the space of $k\times k$ orthogonal matrices. We formalize the notion of information switching through the following definition:
@@ 2.3 Identification properties
Since $\text{rank}(G_{s})=k$, then $\text{rank}(G)=S_{n}k$, and $\text{rank}(G)=\text{rank}(G|\mathbf{w})$ by construction. Thus, since $k<p(p+1)/2$ by assumption, the solution is unique by the Rouché–Capelli Theorem. The uniqueness of the solution for the elements $\psi_{s}^{*}$ ensures that it is impossible to find a representation where for one $h$, $\tilde{\Psi}_{h}$ is a vector of one. ∎
The proof of Theorem 1 is based on proving the existence of unique vectors $\psi_{s}^{*}$ for the specific factors. This guarantees that no alternative representation exists in which, for some $h$, the corresponding $\tilde{\Psi}_{h}$ is a vector of ones. Importantly, it also establishes the uniqueness of the sparsity patterns encoded by the $\Psi$ matrix. That is, any alternative representation inducing the same covariance structure must preserve the subject-specific activation patterns of the specific components. The following corollary formalizes this consequence.
@@
Based on the identifiability findings of Theorem 1, the practical conditions that ensure resistance to information switching in APAFA are discussed in Lemmas 1 and 2.
@@
Under any continuous prior for the elements of $\Gamma$ and truncating its number of columns to K=p(p+1)/21K=p(p+1)/2-1, the information switching for model (7) has zero probability.
@@
From Markov’s inequality $\text{pr}(k>2(p+1)/2)\leq\frac{\alpha^{\varphi}}{p(p+1)/2}\leq\varepsilon$. ∎
A key advantage of our model is that the requirement of $k<p(p+1)/2$ is minimal, as it is customary in factor analysis to have latent factors of a considerably lower dimension than the observed data ($k<p$). Notably, Chandra et al. (2025), in different settings, necessitate the dimension of the specific part to be smaller than that of the shared part implicitly considering the study-specific part as a noise or disturbance parameter. Although we are in different settings, where the number and composition of the groups is not pre-specified, we do not restrict $k$ to be also smaller than $d$. This is crucial in many contexts. For instance, in cancer studies, it is possible to study very different types of cancer simultaneously, identifying a range of specific factors that describe the specific characteristics of each type of cancer, while also uncovering a few common factors that are shared across cancers, regardless of the organ or tissue affected, possibly pertaining to fundamental mechanisms of oncogenesis.
@@
We begin by fixing the ordering of the columns of $\Gamma$. Under this fixed ordering, we will show that the only admissible orthogonal matrices $P\in\mathcal{O}_{k}$ are diagonal with entries $\pm 1$. Moreover, applying any permutation to the columns of $\Gamma$ initially and repeating the argument would yield the same conclusion, i.e. the only admissible $P$ relative to the permuted ordering are diagonal with entries $\pm 1$. Consequently, the general solution requires $P$ to be a signed permutation matrix.
@@
and define $\Gamma_{s}\in\mathbb{R}^{p\times k_{s}}$ to be the submatrix of $\Gamma$ consisting of the columns indexed by $A_{s}$. Notably, these are identifiable thanks to Corollary 1.
The proof proceeds considering that $P_{s}\in\mathcal{O}_{k_{s}}$, leaves $\Gamma^{\prime}_{s}\Gamma^{\prime}_{s}{}^{\top}$ invariant and hence that the marginal distribution of group $s$ is also invariant under such rotations. Given these $P_{s}\in\mathcal{O}_{k_{s}}$ for each $s=1,\dots,S_{n}$, we investigate whether a “global” orthogonal matrix $P\in\mathcal{O}_{k}$ can be constructed by assembling the blocks $P_{s}$. This requires consistency across overlapping blocks, ensuring that each $P_{s}$ remains orthogonal, for example satisfying $\sum_{l\in A_{s}}\|P_{s,hl}\|=1$ for all $h$ and $s$. We will argue in the following that, under the Non-replicable Sparsity Pattern Condition of Definition 2, this is possible only if $P$ is a diagonal matrix with entries $\pm 1$. For any $h=1,\ldots,k$, there exist at least one group loading factor $h$, i.e. $\psi^{*}_{sh}=1$ for at least one $s$. Then we have two possibilities: $i)$ $\psi_{s^{\prime}h}^{*}=0$ for all other $s^{\prime}\neq s$, or ii)ii) $\psi_{s^{\prime}h}^{*}=1$ for at least another $s^{\prime}\neq s$. In the first case $i)$, $h$ is a factor belonging exclusively to group $s$. Clearly, by Definition 2 it is the only exclusive factor. If this is the only factor of group $s$, then the rotation matrix $P_{s}$ reduces to a scalar. Since $P$ is orthogonal, $\|P_{h}\|^{2}=1$ implies $P_{hh}=\pm 1$ and all other entries in the $h$th row/column of $P$ must vanish. If $h$ is not the only factor of group $s$, then $A_{s}$ contains at least another $h^{\prime}$ that is shared among $s$ and at least another $s^{\prime}$. In this case, $h^{\prime}$ is in case ii)ii). Consider then case ii)ii). In order to be orthogonal, matrix $P$ needs to satisfy
 $1=\|P_{h}\|^{2}=\sum_{l=1}^{k}P_{lh}^{2}=\sum_{l\in A_{s^{\prime}}\cap A_{s}}P_{lh}^{2}+\sum_{l\in A_{s^{\prime}}\setminus A_{s}}P_{lh}^{2}+\sum_{l\in A_{s}\setminus A_{s^{\prime}}}P_{lh}^{2}+\sum_{l\notin A_{s}\cup A_{s^{\prime}}}P_{lh}^{2}.$ (8)
@@
 $1=\|P_{{s},h}\|^{2}=\sum_{l\in A_{s}\cap A_{s^{\prime}}}P_{lh}^{2}+\sum_{l\in A_{s}\setminus A_{s^{\prime}}}P_{lh}^{2}$ (9)